# Off-Road Navigation via Implicit Neural Representation of Terrain Traversability

### v2

@@ IV Field Experiments
TABLE IV: Hardware Short Path Navigation Comparison Results
@@
IV-A Short Path Navigation
We chose five sets of experiments to test the capability of all methods in scenarios that present different challenges. Each set of experiments consists of three trials per method. A trial is considered failed if the vehicle ends up in an unrecoverable configuration (e.g. in collision with a tree trunk), in which case we manually terminated the trial. We set the planning horizon of both baseline approaches to $5.0$\mathrm{s}$\$ (used in [frey2024roadrunner] for larger scale experiments) and the number of MPPI samples to 4096. Quantitative results are recorded in Table IV and visualizations are shown in Figure 8. Experiments I and IV are challenging due to tree trunks and roots, Experiment II involves tall vegetation, and Experiments III and V combine both scenarios.
@@ IV-A Short Path Navigation
IV-B Long Path Navigation
We also conducted a long path navigation experiment with TRAIL; the overall route is illustrated in Figure 7. A total of eight interventions were required: five were caused by the platform’s limited traction on wet ground (training data was collected on a sunny day with dry ground), and three were caused by planned trajectories that would have resulted in collisions. Of these three interventions, two were due to small trees that were outside the camera’s field of view and within LiDAR blind spots created by the sensor-stack pillars. The remaining intervention was caused by the trajectory optimizer overly smoothing a pair of consecutive tight turns that were intended to avoid two nearby trees.
IV-C Limitations
@@
V Conclusion
In this work, we proposed and evaluated an off-road navigation framework. The framework consists of a perception module that learns an implicit representation of terrain properties and a planner that utilizes the gradient information from the perception module to perform gradient-based trajectory optimization. The proposed approach is able to navigate effectively in challenging environments. Although we mainly focus on off-road navigation tasks in this paper, the trajectory optimization framework that leverages an implicit neural representation can be applied to a broad class of navigation problems. One future direction is temporal fusion of the traversability prediction to enable longer range navigation and to overcome sensor occlusion.

### v3

@@ V Field Experiments
TABLE V: Hardware Short Path Navigation Comparison Results the Euclidean distances between start and goal are about $9\text{\,}\mathrm{m}$)
@@
V-A Short Path Navigation
We chose five sets of experiments to test the capability of all methods in scenarios that present different challenges. Each set of experiments consists of three trials per method. A trial is considered failed if the vehicle ends up in an unrecoverable configuration (e.g. in collision with a tree trunk), in which case we manually terminated the trial. We set the planning horizon of both baseline approaches to $5.0$\mathrm{s}$\$ (used in [15] for larger scale experiments) and the number of MPPI samples to 4096. Quantitative results are recorded in Table V and visualizations are shown in Figure 6. Experiments I and IV are challenging due to tree trunks and roots, Experiment II involves tall vegetation, and Experiments III and V combine both scenarios.
@@ V-A Short Path Navigation
V-B Long Path Navigation
We also conducted a long-range navigation experiment with TRAIL (route shown in Figure 5). Eight interventions were required. Five interventions were caused by limited traction on wet ground, a hardware-limited regime in which the platform’s stock tires could not provide sufficient friction for wet off-road terrain, leading to wheel slip even under conservative commands. The remaining three interventions were safety-related: two were due to small trees outside the camera’s field of view and within LiDAR blind spots introduced by the sensor-stack pillars, and one was caused by the trajectory optimizer overly smoothing consecutive tight turns intended to pass between nearby trees. Potential mitigations include equipping tires suitable for wet/off-road conditions, and algorithmically, re-normalizing the bumpiness signal with a short calibration run on the target surface and including wet/low-traction conditions during data collection.
V-C Limitations
@@
VI Conclusion
In this work, we proposed and evaluated an off-road navigation framework. The framework consists of a perception module that learns an implicit representation of terrain properties and a planner that utilizes the gradient information from the perception module to perform gradient-based trajectory optimization. The proposed approach is able to navigate effectively in challenging environments. One future direction is temporal fusion of the traversability prediction to enable longer range navigation and to overcome sensor occlusion.